# Hidden in the Noise: Unveiling Backdoors in Audio LLMs Alignment through Latent Acoustic Pattern Triggers

### v1

@@ Introduction
[…]
@@
[…]
@@
  • Loss Differential Variance ($\text{Var}(\nabla\mathcal{L})$): Measures the spread of $\nabla\mathcal{L}$ values around their mean, indicating the consistency of the deviation. For a set of $N$$\mathcal{L}$ values, its variance is calculated as $\text{Var}(\nabla\mathcal{L})=\frac{1}{N}\sum_{i=1}^{N}(\nabla\mathcal{L}_{i}-\overline{\nabla\mathcal{L}})^{2}$, where $\overline{\nabla\mathcal{L}}$ represents the mean of all $\nabla\mathcal{L}$ values.
@@
As illustrated in Figure 3, the loss trends for ALLMs trained with acoustic triggers closely mirror those trained on clean samples, visually confirming the covertness of our backdoor injections. This visual consistency aligns with quantitative findings, as shown by the consistently low $\nabla\mathcal{L}$) and, for the majority of data, generally low $\nabla\mathcal{L}$)||| values presented in Table 1. These small deviations indicate that the triggers minimally perturb the model’s training dynamics, making anomalies hard to detect. Notably, the frequent occurrence of negative $\nabla\mathcal{L}$) values suggests that triggered sample losses are often even lower than clean losses, further enhancing stealth by avoiding the typical loss increase. Overall, Qwen2.5-Omni (Fan et al. 2025) consistently demonstrates superior covertness, with its loss profile being the least impacted, confirming the feasibility of embedding hidden backdoors in ALLMs without discernible disruption to their learning.

### v2

@@ Introduction
in which $\phi_{e}$ transforms audio tokens to embeddings while $\phi_{t}$ converts text tokens to the same embedding space. The symbol $\|$ denotes concatenation.
@@
    $A_{\text{trigger}}(t)=\alpha\cdot A(t)=\alpha\cdot\int_{-\infty}^{t}h(t-\tau)\cdot s(\tau)\,d\tau,$ (9)
@@
  • Loss Differential Variance ($\text{Var}(\nabla\mathcal{L})$): Measures the spread of $\nabla\mathcal{L}$ values around their mean, indicating the consistency of the deviation. For a set of NN $\nabla\mathcal{L}$ values, its variance is calculated as $\text{Var}(\nabla\mathcal{L})=\frac{1}{N}\sum_{i=1}^{N}(\nabla\mathcal{L}_{i}-\overline{\nabla\mathcal{L}})^{2}$, where $\overline{\nabla\mathcal{L}}$ represents the mean of all $\nabla\mathcal{L}$ values.
@@
As illustrated in Figure 3, the loss trends for ALLMs trained with acoustic triggers closely mirror those trained on clean samples, visually confirming the covertness of our backdoor injections. This visual consistency aligns with quantitative findings, as shown by the consistently low Var($\nabla\mathcal{L}$) and, for the majority of data, generally low ||CV($\nabla\mathcal{L}$)$|$ values presented in Table 1. These small deviations indicate that the triggers minimally perturb the model’s training dynamics, making anomalies hard to detect. Notably, the frequent occurrence of negative CV($\nabla\mathcal{L}$) values suggests that triggered sample losses are often even lower than clean losses, further enhancing stealth by avoiding the typical loss increase. Overall, Qwen2.5-Omni (Fan et al. 2025) consistently demonstrates superior covertness, with its loss profile being the least impacted, confirming the feasibility of embedding hidden backdoors in ALLMs without discernible disruption to their learning.